CRY+VIS Interface Meeting
Date: 2022/4/15 13:00-13:50
Participants: Aso, Ushiba, Tomaru, Akutsu, Kimura, Ikeda, Miyoki, Takahashi, Sato, Hirata
- IX: F0 overload
- IY: Health check. Go to finalization (WAB, Frame adjustment).
- EY: TM actuator problem (flip cable, regluing of magnet).
- Type-B/Bp/C: BFH LVDT was recovered in PRM. Power supply trouble was due to the broken cooling fan. Leak in the IMC chamber was recovered.
- Replacement of flip cables: Cables for TMS were ordered. Feedthrough flange was replaced with the blank flange in OMC.
- TM height setting: ITMY.
- Oplev: Calibration results were summarized. Windshield in the next.
- ACC: Investigating the noise source.
- Commissioning in air:
- Commissioning in vacuum: Alignment of the MCE and the X-arm is scheduled in the next week.
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers:
- Heater control: Test is pending.
- TMS: Pressure fluctuation was due to the valve operations. Two vacuum gauges were broken.
Date: 2022/4/15 14:00-14:30
Participants: Takahashi, Ushiba, Fabian, Washimi, Tanaka, Uchiyama
- Recovery of the TM actuator in EY.
- Saturation of the F0 keystone height.
- We replaced a broken LVDT in PRM. Namely, the one for the BF Damper horizontal LVDTs:
- In the old board, we measured the resistance values determining two gains, and copied them into the new board.
- After installing the new board, we adjusted the phases in order to achieve maximum output.
- Then, in order to calculate the new LVDT calibration factors, we measured the transfer functions from Y actuation to H1, H2 and H3 outputs, and compared with previous measurements.
- The DC amplitudes and phases of transfer functions are the same as before.
- Hirata-san continued working in the design of the Traverser limit switches. Differences between the real hardware and the 3D-CAD have made the process difficult, but he's now working on a final version. He has reported his work at the ATC-KAGRA meeting.
- Takahashi-san realized that the LVDT drivers used in PR3 and PRM are using a reference signal with an amplitude of 12 Vpp.
- Replaced the feedthrough flange with the blank flange in OMC.
- Type-A team: Finalization in IY.
- Type-B team: SF FR grounded problem in SR3.
(Date in Kamioka)
The week from 4/18
- Takahashi: 4/19-22
- Sato: 4/19-22
- Hirata: 4/19-22
The week from 4/25
- Takahashi: 4/25-28
- Sato: 4/25-28
- Hirata: 4/25-28